OpenStudy (anonymous):

Can someone please help me with this...... If y=x(lnx)^2 then (dy/dx)=

OpenStudy (anonymous):

Product rule. d(uv)=u'v+uv'. Also, don't forget to use the chain rule when calculating the derivative of (ln x)^2.

OpenStudy (anonymous):

\[\begin{align*}y&=x(\ln x)^2\\ \frac{dy}{dx}&=\frac{d}{dx}\left[x(\ln x)^2\right]\\ &=\frac{d}{dx}[x]\cdot(\ln x)^2 + x\cdot\frac{d}{dx}\left[(\ln x)^2\right] \end{align*}\] Applying power rule and chain rule on the right-most term, you have \[\frac{d}{dx}\left[(\ln x)^2\right]=2\ln x\cdot \frac{d}{dx}[\ln x]\]

OpenStudy (anonymous):

would the answer be (lnx)(2+xlnx)

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

It's close, but no.

OpenStudy (anonymous):

what is wrong

OpenStudy (anonymous):

You shouldn't have a power of \(x\) in your final answer. Remember, you have \[\color{red}{x}\cdot2\ln x\cdot\color{blue}{\frac{d}{dx}\left[\ln x\right]}\] What can you tell about the highlighted terms?

OpenStudy (anonymous):

(lnx)(2+lnx)

OpenStudy (anonymous):

Yup, that's it.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

You're welcome